OpenStudy (anonymous):

I GIVE MEDALS. A cyclist rides in a circle with speed 3.2 m/s. What is his centripetal acceleration if the circle has a radius of 12 m? A. 1.4 m/s^2 B. 0.85 m/s^2 C. 0.27 m/s^2 D. 3.8 m/s^2

OpenStudy (anonymous):

\[a = \frac{v^2}{r} = \frac{(3.2 \text{ ms}^{-1})^2}{12\text{ m}}\]

OpenStudy (anonymous):

I got 0.27 (c) is that correct? @broken_symmetry

OpenStudy (anonymous):

That's just \(\frac{3.2}{12}\). You didn't square the speed.

OpenStudy (anonymous):

so it would be .85? @broken_symmetry

OpenStudy (anonymous):

Yep